# L4DR: LiDAR-4DRadar Fusion for Weather-Robust 3D Object Detection

### v1

@@ Introduction
Implement Details.
We implement L4DR with PointPillars (Lang et al. 2019), the most commonly used base architecture in radar-based, LiDAR and 4D radar fusion-based 3D object detection. This can effectively verify the effectiveness of our L4DR and avoid unfair comparisons caused by inherent improvements in the base architecture. We set $\tau$ in section 3.2 as 0.3 while training and 0.2 while inferring (more discussions can be found in supplementary material A.2.). We conduct all experiments with a batch size of 16 on 2 RTX 3090 GPUs. Other parameter settings refer to the default official configuration in the OpenPCDet (Team et al. 2020) tool.
Dataset and Evaluation Metrics
K-Radar dataset.
The K-Radar dataset (Paek, Kong, and Wijaya 2022) contains 58 sequences with 34944 frames of 64-line LiDAR, camera, and 4D radar data in various weather conditions. According to the official K-Radar split, we used 17458 frames for training and 17536 frames for testing. We adopt two evaluation metrics for 3D object detection: $AP_{3D}$ and $AP_{BEV}$ of the class ”Sedan” at IoU = 0.5. We also provide more quantitative results for other IoU thresholds as well as using the latest v2.1 version of the label (see supplementary material).
View-of-Delft (VoD) dataset.
The VoD dataset (Palffy et al. 2022) contains 8693 frames of 64-line LiDAR, camera, and 4D radar data. Following the official partition, we divided the dataset into a training and validation set with 5139 and 1296 frames. Meanwhile, to explore the performance under different fog intensities, following a series of previous work (Qian et al. 2021; Li et al. 2022), we similarly performed fog simulations (Hahner et al. 2021) (with fog level $\mathcal{L}$ from 0 to 4, fog density ($\alpha$) = [0.00, 0.03, 0.06, 0.10, 0.20]) on the VoD dataset and 4D Radar remains unchanged to simulate weather robustness. We express this as the Vod-Fog dataset in the following. Noteworthy, we used two evaluation metrics groups. The VoD official metrics are used to compare better the results reported by previous state-of-the-art methods. The KITTI official metrics are for better demonstrating and analyzing the performance of ”easy”, ”moderate” and ”hard” objects with different difficulties under foggy weather, for more details of the metrics see the supplementary material.
Results on K-Radar Adverse Weather Dataset
Following 3D-LRF (Chae, Kim, and Yoon 2024), we compare our L4DR with LiDAR only, 4D radar only, and LiDAR-4D radar fusion-based 3D object detection methods: PointPillars (Lang et al. 2019), RTNH (Paek, KONG, and Wijaya 2022), InterFusion (Wang et al. 2022b) and 3D-LRF (Chae, Kim, and Yoon 2024). The results in Table 2 highlight the superior performance of our L4DR model in all weather conditions. Our L4DR model surpasses 3D-LRF by 8.3% in total $AP_{3D}$. This demonstrates that compared to previous fusion frameworks, our proposed fusion method utilizes the advantages of LiDAR and 4D Radar more effectively. Note that we only compare 3D-LRF (Chae, Kim, and Yoon 2024) on the K-Radar dataset because its code is not open-sourced, the results of 3D-LRF are available on the K-Radar dataset only.
Meanwhile, it is worth noting that the performance in many bad weather conditions instead significantly exceeds the performance in normal weather (e.g., Overcast, Fog, etc.). This is a phenomenon also reflected in the official benchmark of the K-Radar dataset, and we discuss this counter-intuitive phenomenon in detail in the supplementary material. We also discuss other valuable results such as different IoU thresholds and new version labels.
@@
We systematically evaluated each component, with the results summarized in Table 5. The $1^{st}$ row represents the performance of the LiDAR-only baseline model. Subsequent enhancements were observed with the addition of MME, as depicted in the $2^{nd}$ row, showcasing significant performance gains attributed to the radar. This performance improvement becomes more significant with increasing fog levels, highlighting the weather resilience of the 4D radar. The $3^{rd}$ row demonstrates further optimization of data fusion between the modalities through noise filtering in the 4D radar using FAD. However, the $4^{th}$ row indicates that the performance boost from incorporating the {IM}2 model alone was not substantial, primarily due to feature redundancy introduced by the {IM}2 backbone. This issue was effectively addressed by utilizing the MSGF module in the $5^{th}$ row, leading to the most optimal performance.
@@
MME {IM}2 FAD MSGF W/o Fog $\mathcal{L}$ = 1 $\mathcal{L}$ = 2 $\mathcal{L}$ = 3 $\mathcal{L}$ = 4
@@
Table 5: Effect of each component proposed in L4DR, tested on Vod-Fog split. 3D mAP represents the average 3D AP of all classes in Mod. difficulty. $\mathcal{L}$ denotes fog level. The first line does not use any modules indicating PointPillar model that only uses LiDAR.
Comparison with other multi-modal feature fusion.
We compared different multi-modal feature fusion blocks, including basic concatenation (Concat.) and various attention-based methods such as Transformer-based (Vaswani et al. 2017) Cross-Modal Attention (Cross-Attn.) and Self-Attention (Self-Attn.), SE Block (Hu, Shen, and Sun 2018), and CBAM Block (Woo et al. 2018), see supplementary material for detailed implement. Experimental results (Table 6) show that while attention mechanisms outperform concatenation to some extent, they do not effectively address the challenge of fluctuating features under varying weather conditions. In contrast, our proposed MSGF method, focusing on significant features of LiDAR and 4D radar, achieves superior performance and robustness under different weather.
@@ Conclusion
In this paper, we analyzed the challenges of fusing LiDAR and 4D radar in adverse weather and proposed L4DR, an effective LiDAR and 4D radar fusion method. We provide an innovative and feasible solution for achieving weather-robust outdoor 3D object detection in various weather conditions. Our experiments on VoD and K-Radar datasets have demonstrated the effectiveness and superiority of our method in various simulated fog levels and real-world adverse weather compared to other fusion methods, as well as other modalities. In summary, our proposed L4DR fusion method not only offers a promising solution for robust outdoor 3D object detection in adverse weather conditions but also sets a new benchmark for performance and robustness compared to existing fusion techniques, paving the way for enhanced safety and reliability in autonomous driving and other applications.
Limitations. While the introduction of the {IM}2 and MSGF modules has allowed the model to focus on more salient features, it inevitably introduces additional computations that reduces the computational efficiency to a certain extent. The inference speed is reduced to about 10 FPS, which is just enough to satisfy the real-time threshold (equal to the LiDAR acquisition frequency), but the computational performance optimization of our L4DR method is a valuable future area of research.
@@
For the VoD dataset, following KITTI (Geiger, Lenz, and Urtasun 2012), we calculate the 3D Average Precision (3D AP) across 40 recall thresholds (R40) for different classes. Also, following VoD’s (Palffy et al. 2022) evaluation metrics, we calculate class-wise AP and mAP averaged over classes. The calculation encompasses the entire annotated region (camera FoV up to 50 meters) and the ”Driving Corridor” region ([-4 m ¡ x ¡ +4 m, z ¡ 25 m]). For both KITTI metrics and VoD metrics, for AP calculations, we used an IoU threshold specified in VoD, requiring a 50% overlap for car class and 25% overlap for pedestrian and cyclist classes.
Experimental Visualization Results
Experiments of the hyperparameter $\lambda$ in FAD
Foreground Semantic Segmentation Results in FAD
$\lambda$ in testing Recall IoU PA
0.1 84.35 38.07 88.43
0.2 78.04 50.11 93.45
0.3 73.15 54.14 94.78
0.5 65.52 54.39 95.52
Table 8: Experimental Recall, IoU, and PA performance results of foreground semantic segmentation in FAD ($\lambda$ in training is 0.3).
We also tested the experimental results of FAD for the denoising stage in semantic segmentation. We used Recall, IoU, and Point Accuracy (PA) as the evaluation metrics, as shown in Table 8. With the increase of hyperparameter $\lambda$, Recall decreases and IoU and PA gradually increase. At $\lambda$ = 0.5, we obtain the best IoU and PA, but the lowest recall. It verifies the correctness of our denoising algorithm for semantic segmentation. However, we find that the performance of the 3D object detection task is not as good when $\lambda$ is higher because losing more foreground points is more detrimental to object detection than retaining some background points.

### v3

@@ Introduction
Implement Details.
We implement L4DR with PointPillars (Lang et al. 2019), the most commonly used base architecture in radar-based, LiDAR and 4D radar fusion-based 3D object detection. This can effectively verify the effectiveness of our L4DR and avoid unfair comparisons caused by inherent improvements in the base architecture. We set $\tau$ in section 3.2 as 0.3 while training and 0.2 while inferring. We conduct all experiments with a batch size of 16 on 2 RTX 3090 GPUs. Other parameter settings refer to the default official configuration in the OpenPCDet (Team et al. 2020) tool.
Dataset and Evaluation Metrics
K-Radar dataset.
The K-Radar dataset (Paek, Kong, and Wijaya 2022) contains 58 sequences with 34944 frames of 64-line LiDAR, camera, and 4D radar data in various weather conditions. According to the official K-Radar split, we used 17458 frames for training and 17536 frames for testing. We adopt two evaluation metrics for 3D object detection: $AP_{3D}$ and $AP_{BEV}$ of the class ”Sedan” at IoU = 0.5.
View-of-Delft (VoD) dataset.
The VoD dataset (Palffy et al. 2022) contains 8693 frames of 64-line LiDAR, camera, and 4D radar data. Following the official partition, we divided the dataset into a training and validation set with 5139 and 1296 frames. All of the methods used the official radar with 5 scans accumulation and single frame LiDAR. Meanwhile, to explore the performance under different fog intensities, following a series of previous work (Qian et al. 2021; Li et al. 2022), we similarly performed fog simulations (Hahner et al. 2021) (with fog level $\mathcal{L}$ from 0 to 4, fog density ($\alpha$) = [0.00, 0.03, 0.06, 0.10, 0.20]) on the VoD dataset and kept 4D Radar unchanged to simulate various weather conditions. We named it the Vod-Fog dataset in the following. Noteworthy, we used two evaluation metrics groups on the VoD dataset. The VoD official metrics are used to compare the results reported by previous state-of-the-art methods. The KITTI official metrics are used to demonstrate and analyze the performance of ”easy”, ”moderate” and ”hard” objects with different difficulties under foggy weather.
Results on K-Radar Adverse Weather Dataset
Following 3D-LRF (Chae, Kim, and Yoon 2024), we compare our L4DR with different modality-based 3D object detection methods: PointPillars (Lang et al. 2019), RTNH (Paek, KONG, and Wijaya 2022), InterFusion (Wang et al. 2022b) and 3D-LRF (Chae, Kim, and Yoon 2024). The results in Table 1 highlight the superior performance of our L4DR model on the K-Radar dataset. Our L4DR model surpasses 3D-LRF by 8.3% in total $AP_{3D}$. This demonstrates that compared to previous fusion, our method utilizes the advantages of LiDAR and 4D Radar more effectively. We observed that all methods perform better in many adverse weather conditions (e.g., Overcast, Fog, etc.) than in normal weather. A possible reason is the distribution differences of labeled objects in different weather conditions. We discuss this counter-intuitive phenomenon in detail in the supplementary material, as well as other valuable results, such as different IoU thresholds and new version labels, for a more comprehensive comparison. Note that we only compare 3D-LRF on the K-Radar dataset because its code is not open-sourced, the results of 3D-LRF are available on the K-Radar dataset only.
@@
We systematically evaluated each component, with the results summarized in Table 4. The $1^{st}$ row represents the performance of the LiDAR-only baseline model. Subsequent $2^{nd}$ row and $3^{rd}$ row are fused by directly concatenating the BEV features from LiDAR and 4D radar modality. The results showing the enhancements were observed with the addition of MME and FAD respectively, highlighting that our fusion method fully utilizes the weather robustness of the 4D radar while excellently handling the noise problem of the 4D radar. The $4^{th}$ row indicates that the performance boost from incorporating the {IM}2 model alone was not substantial, primarily due to feature redundancy introduced by the {IM}2 backbone. This issue was effectively addressed by utilizing the MSGF module in the $5^{th}$ row, leading to the most optimal performance.
@@
MME FAD {IM}2 MSGF W/o Fog $\mathcal{L}$ = 1 $\mathcal{L}$ = 2 $\mathcal{L}$ = 3 $\mathcal{L}$ = 4
@@
Table 4: Effect of each component proposed in L4DR, tested on Vod-Fog split. 3D mAP represents the average 3D AP of all classes in Mod. difficulty. $\mathcal{L}$ denotes fog level.
Comparison with other feature fusion.
We compared different multi-modal feature fusion blocks, including basic concatenation (Concat.) and various attention-based methods such as Transformer-based (Vaswani et al. 2017) Cross-Modal Attention (Cross-Attn.) and Self-Attention (Self-Attn.), SE Block (Hu, Shen, and Sun 2018), and CBAM Block (Woo et al. 2018), see supplementary material for detailed fusion implements. Experimental results (Table 5) show that while attention mechanisms outperform concatenation to some extent, they do not effectively address the challenge of fluctuating features under varying weather conditions. In contrast, our proposed MSGF method, focusing on significant features of LiDAR and 4D radar, achieves superior performance and robustness under different weather.
@@
Foreground semantic segmentation results in FAD
To explore the 4D radar denoising ability of FAD, we tested different $\lambda$ in the inference stage, as shown in Table 6. With hyperparameter $\lambda$ increase, recall gradually decreases, and denoise rate, mIoU, and PA gradually increase. We find that the performance of the 3D object detection is not excellent when $\lambda$ is relatively high (0.5). The reason is missing foreground points is more unacceptable than retaining some noise for object detection. Additionally, we also discuss the effect of $\lambda$ on the detection performance in the supplementary material.
$\lambda$ in testing Denoise Rate Recall mIoU PA
0.1 88.76 84.35 38.07 88.43
0.2 94.68 78.04 50.11 93.45
0.3 96.51 73.15 54.14 94.78
0.5 97.92 65.52 54.39 95.52
Table 6: Denoise rate, recall, mIoU, and PA (Point Accuracy) of foreground segmentation in FAD ($\lambda$ in training is 0.3).
@@ Conclusion
In this paper, we analyzed the challenges of fusing LiDAR and 4D radar in adverse weather and proposed L4DR, an effective LiDAR and 4D radar fusion method. We provide an innovative and feasible solution for achieving weather-robust outdoor 3D object detection in various weather conditions. Our experiments on VoD and K-Radar datasets have demonstrated the effectiveness and superiority of our method in various simulated fog levels and real-world adverse weather. In summary, our proposed L4DR fusion method not only offers a promising solution for robust outdoor 3D object detection in adverse weather conditions but also sets a new benchmark for performance and robustness compared to existing fusion techniques, paving the way for enhanced safety and reliability in autonomous driving and other applications.
Limitations. While the introduction of the {IM}2 and MSGF modules has allowed the model to focus on more salient features, it inevitably introduces additional computations that reduce the computational efficiency to a certain extent. The inference speed is reduced to about 10 FPS, which is just enough to satisfy the real-time threshold (equal to the LiDAR acquisition frequency), but the computational performance optimization is a valuable future area of research.
@@
Discussion about the Counter-Intuitive Results on K-Radar dataset
As shown by the results in the main paper: K-Radar inverse weather dataset experiment, all methods perform better in many severe weather conditions (e.g., cloudy, foggy, etc.) than in normal weather. After analyzing the data and experimental results, a possible main reason is the distribution differences of labeled objects in different weather conditions.We will verify this in two respects.
The performance of 4D Radar-based method in different weather condition.
Benefiting from the sensor weather robustness of 4D radar, the 4D Radar-based method receives little performance impact from weather. Therefore, the performance of the 4D Radar-based method under each weather can be used to determine the difference in difficulty caused by the distribution of labeled objects for different weather on the K-radar dataset. As shown in the results demonstrated by the 4D Radar-based method (RTNH, PointPillars) in Tables 8 and 9, it can be found that the performance of the 4D Radar-based method is significantly worse than that of most of the bad weather in normal weather. This result illustrates that labeled objects in bad weather scenarios are much easier to detect in the K-Radar dataset. This makes it possible that even when LiDAR is degraded in bad weather, the performance is still higher than in normal weather due to the fact that the labeled targets in the scene are easier to detect.
Figure 8: Distribution of labeled bounding boxes in normal (left) and severe (right) weather in the K-Radar dataset.
Labeled Object Distance-Number Distribution of K-Radar Dataset.
We counted the number of labeled objects at different distances by weather category (normal, severe), as shown in Figure 8. It can be found that under normal weather, the proportion of the number of labeled objects (26.80%) at long distances (50-75m) is greater than that under severe weather (22.66%), while the proportion of the number of labeled objects (36.78%) at closed distances (0-25m) is smaller than that under severe weather (40.0%). This data reflects the difficulty of the scenarios in different weather, explaining the counter-intuitive results on K-Radar dataset.
Experiments of the Hyperparameter $\lambda$ in FAD
@@
For the VoD dataset, following KITTI (kitti), we calculate the 3D Average Precision (3D AP) across 40 recall thresholds (R40) for different classes. Also, following VoD’s (Palffy et al. 2022) evaluation metrics, we calculate class-wise AP and mAP averaged over classes. The calculation encompasses the entire annotated region (camera FoV up to 50 meters) and the ”Driving Corridor” region ([-4 m ¡ x ¡ +4 m, z ¡ 25 m]). For both KITTI metrics and VoD metrics, for AP calculations, we used an IoU threshold specified in VoD, requiring a 50% overlap for car class and 25% overlap for pedestrian and cyclist classes.